OpenStudy (anonymous):

find the velocity and acceleration of the process: x=10e-2t.cos (5t + π/2)

OpenStudy (anonymous):

velocity is first derivative, acceleration the second derivative

OpenStudy (agreene):

^ assuming this is a displacement vector.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

pls help solve the first and second derivative

OpenStudy (anonymous):

thanks @ satellite. I have class now. take care

OpenStudy (anonymous):

@satellite73

OpenStudy (agreene):

\[x'= \frac{1}{10}(-5 t \cos(5 t) - \sin(5 t))\] \[x''= \frac52 t \sin(5 t)-\cos(5 t)\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

x=t.e-t2/2

OpenStudy (anonymous):